The parks departement plans to fence in a 2304 sq meter rectangular area park. What should be the dimensions in order to use the least amount of fencing.
Answer by Cintchr(481)

Area = 2304 Perimeter = ???
A=1 * 2304 and P=2(1+2304)= 4610
A=2 * 1152 and P=2(2+1152)= 2308
A=3 * 768 and P=2(3+768)= 1542
A=4 * 576 and P=2(4+576)= 1160
If we continue this pattern we get to ....
A=48 * 48 and P=2(48+48)= 192
The smallest amout of fencing with the stated required area.
whenever you have a problem like this .... look for the shape to be a perfect square.